OpenStudy (anonymous):

At a factory that produces pistons for cars, Machine 1 produced 801 satisfactory pistons and 89 unsatisfactory pistons today. Machine 2 produced 783 satisfactory pistons and 117 unsatisfactory pistons today. Suppose that one piston from Machine 1 and one piston from Machine 2 are chosen at random from today's batch. What is the probability that the piston chosen from Machine 1 is unsatisfactory and the piston chosen from Machine 2 is satisfactory ?

OpenStudy (anonymous):

Total pistons of A=890 P(unsatisfactory pistons)=89/890=0.1 Total pistons of B=783+117=900 P(satisfactory pistons)=783/900=0.87 Final ans would be 0.1x0.87=0.0087 (87/1000)